OpenStudy (anonymous):

Before leaving the ground, an aircraft moves with constant acceleration and travels 720m in 12s from rest. It then leaves the ground. Determine (a) the acceleration (b) the speed with which it leaves the ground, (c) the distance covered during the first second and during the twelfth second. answers (10 ms^-2, 120 ms^-1, 5m, 115m)

OpenStudy (anonymous):

Speed at which it leaves the ground is derived by \[Speed=Distance \div time\].Therefore,speed=720/12=60

OpenStudy (anonymous):

first second will be distance= \[=speed \times time\]....equal to 60 times 1 = 60

OpenStudy (anonymous):

for 12th second..distance=\[60\times12\]=720m

OpenStudy (anonymous):

Acceleration = 0 since it is constant